OpenStudy (anonymous):

Please solve. (Algebra 2)

OpenStudy (anonymous):

\[\ln x - \ln 5 = -1\]

OpenStudy (anonymous):

@thomaster ?

OpenStudy (whpalmer4):

\[\ln x - \ln 5 = -1\] Do you remember the property of logs with division? \[\log a - \log b = \log \frac{a}{b}\] Can you use that to combine the left side of the equation into one logarithm?